Contents
What is the difference between the two probabilities?
If the probability of an event occurring is Y, then the probability of the event not occurring is 1-Y. (Example: If the probability of an event is 0.80 (80%), then the probability that the event will not occur is 1-0.80 = 0.20, or 20%.
How do I find my PA and B?
Formula for the probability of A and B (independent events): p(A and B) = p(A) * p(B). If the probability of one event doesn’t affect the other, you have an independent event. All you do is multiply the probability of one by the probability of another.
How do you add two probabilities?
Just multiply the probability of the first event by the second. For example, if the probability of event A is 2/9 and the probability of event B is 3/9 then the probability of both events happening at the same time is (2/9)*(3/9) = 6/81 = 2/27.
How do you know if probabilities are mutually exclusive?
Two events are mutually exclusive if they cannot occur at the same time. Another word that means mutually exclusive is disjoint. If two events are disjoint, then the probability of them both occurring at the same time is 0.
When can you add two probabilities?
Addition Rule 1: When two events, A and B, are mutually exclusive, the probability that A or B will occur is the sum of the probability of each event. Addition Rule 2: When two events, A and B, are non-mutually exclusive, there is some overlap between these events.
What’s the difference between the probability and the odds?
The probability that an event will occur is the fraction of times you expect to see that event in many trials. Probabilities always range between 0 and 1. The odds are defined as the probability that the event will occur divided by the probability that the event will not occur.
What is the probability of an event not occurring?
If the probability of an event occurring is Y, then the probability of the event not occurring is 1-Y. (Example: If the probability of an event is 0.80 (80%), then the probability that the event will not occur is 1-0.80 = 0.20, or 20%.
What is the probability that a horse will win a race?
If the horse runs 100 races and wins 80, the probability of winning is 80/100 = 0.80 or 80%, and the odds of winning are 80/20 = 4 to 1. NOTE that when the probability is low, the odds and the probability are very similar.
How to calculate the probability of getting a length range?
I would like to compare the probability of getting a particular length range in two different conditions. e.g. P 1 = P ( 0.1 < L < 0.2, f e r t i l i z e r A), P 2 = P ( 0.1 < L < 0.2, f e r t i l i z e r B). So I create 2 histograms for the lengths corresponding to each of the 2 fertilizers, H A and H B.